Apply the second derivative test to find the relative extrema of the function f(x)=ln(x2+x+1)

  1. Apply the second derivative test to find the relative extrema of the function

f(x)=ln(x2+x+1)
Answer #1

© ¢3) = ln (42+x+1) = f(x) = 2x+4 (x?+2+1) For max/min, f(x) = 0 i 2x+1 =0 (72+2+12 = x= -1/2 (critical point) f/ = (3+2+1